### v3

# 1]Universität Hamburg, Fakultät für Mathematik, Informatik und Naturwissenschaften, Fernerkundung & Assimilation, Bundesstr. 53, 20146 Hamburg, Deutschland

Abstract
A hybrid 4D-variational data assimilation method for chaotic climate models is introduced using the Lorenz ’63 model. This approach aims to optimise an Earth system model (ESM), for which no adjoint exists, by utilising an adjoint model of a different, potentially simpler ESM. The technique relies on synchronisation of the model to observed time series data employing the dynamical state and parameter estimation (DSPE) method to stabilise the tangent linear system by reducing all positive Lyapunov exponents to negative values. Therefore, long windows can be used to improve parameter estimation. In this new extension a second layer of synchronisation is added between the two models, with and without an adjoint, to facilitate linearisation around the trajectory of the model without an adjoint. The method is conceptually demonstrated by synchronising two Lorenz ’63 systems, representing two ESMs, one with and the other without an adjoint model. Results are presented for an idealised case of identical, perfect models and for a more realistic case in which they differ from one another. If employed with a coarser ESM with an adjoint, the method will save computational power as only one forward run with the full ESM per iteration needs to be carried out. It is demonstrated that there is negligible error and uncertainty change compared to the ‘traditional’ optimisation of full ESM with an adjoint. In a variation of the method outlined, synchronisation between two identical models can be used to filter noisy data. This reduces optimised parametric model uncertainty by approximately one third. Such a precision gain could prove valuable for seasonal, annual, and decadal predictions.
[…]
The Earth system can be realistically described using numerical models that capture the system’s processes and time evolution. ESMs can be used to forecast future states of the system provided that the past and present are known. Data assimilation is a powerful tool that combines observations and a numerical model representing the dynamic principles governing the system to generate an estimation of its state (Wunsch and Heimbach, 2006; Nichols, 2010) in an attempt to further improve the ESMs predictive skills.
Generally, there are two common approaches to incorporate observations into a model, a sequential data assimilation scheme (Bertino et al., 2003) and the variational approach (Le Dimet and Talagrand, 1986). In this work, we use the variational data assimilation approach, which involves minimising a scalar, also called cost function, defined as the quadratic misfit between the observational and model data within an assimilation time window. This method is commonly known as four-dimensional variational assimilation (4D-Var.) (Rabier and Liu, 2003), which utilises an adjoint of the model to iteratively minimise the model-data misfit, by adjusting control variables (Tett et al., 2017; Lyu et al., 2018; Köhl and Willebrand, 2002; Allaire, 2015; Navon, 2009). Aside from data assimilation purposes, adjoint models have been widely used for sensitivity analysis in meteorology and oceanography (Hall et al., 1982; Hall and Cacuci, 1983; Hall, 1986; Marotzke et al., 1999; Stammer et al., 2016); for instance, calculating sensitivity with respect to lateral boundary condition (Gustafsson et al., 1998), estimating the sensitivity of the 2m surface temperature with respect to the sea surface temperature, sea ice, sea surface salinity (Stammer et al., 2018), and many more.
Due to the non-linearities within ESMs, the relevance of adjoint method, and thus the gradient of the cost function in reducing model-data misfits, can be limited by the predictability time scale. This can lead to spikes in the adjoint sensitivities as a result of multiple local minima in the cost-function. Under such circumstances spikes occur in the estimated gradients and the cost function becomes very rough by showing an increasing number of local minima (Köhl and Willebrand, 2002; Lea et al., 2000). Fortunately, the problem can be mitigated by synchronisation which removes the non-linear or chaotic dynamics from the adjoint model leading to a smooth cost function (Abarbanel et al., 2010; Sugiura et al., 2014). This method allows extension of the assimilation window beyond the predictability time-scale, provided that sufficient observations are available. Another challenge of applying an adjoint model to any state-of-the-art system is that these ESMs have a very large number of state variables $\mathcal{O}(10^{7}-10^{8})$, and the computation of its adjoint would be very time consuming and require significant volumes of memory (Stammer et al., 2018).
To mitigate both problems, we propose a novel framework where we use two climate models both coupled through synchronisation, one with a high resolution and the other with coarse resolution for which an adjoint exists. In this context we only use the adjoint of the low-resolution model to estimate the parameters that optimise the high-resolution model; i.e., that brings the model closer to the assimilated observations. A pre-requisite of this method is that the parameters being optimised must be the same in both models. The objective of this paper is to quantify the precision and the benefit of such a synchronised data assimilation approach. We perform this test conceptually using a Lorenz ’63 model system.
The Lorenz ’63 system (Lorenz, 1963) is a well established proxy model to study chaotic fluid systems, such as the atmosphere (Gauthier, 1992; Miller et al., 1994; Pires et al., 1996; Stensrud and Bao, 1992; Kravtsov and Tsonis, 2021; Huai et al., 2017; Yang et al., 2006; Daron and Stainforth, 2015; Errico, 1997). The advantage is that it can be used to quantitatively evaluate the parameter dependence of the system prior to application in a full model. New modelling techniques can thus be trialled in relatively fast experiments (Pasini and Pelino, 2005; Tandeo et al., 2015; Goodliff et al., 2020; Marzban, 2013; Yin et al., 2014). It can also be used in a wide range of other applications (Du and Shiue, 2021; Cameron and Yang, 2019; Pelino and Maimone, 2007). The system generates a three dimensional, time varying trajectory which with variation of both model parameters and/or initial conditions will produce very different trajectories. Thus, it is an ideal test bed for non-linear modelling in a number of fields (Hirsch et al., 2013). In a previous study (Lyu et al., 2018) using the Lorenz ’63 model, the assimilation primarily focused on the fit of a single parameter $\rho$ and the initial conditions $(x,y,z)$ in a single model with an adjoint. This study will expand on the previous to fit all three model parameters simultaneously and use a model with adjoint to optimise the parameters of a model without one.
The structure of the remaining paper is as follows: In Section 1 we introduce the model, outline the methodology of how these models are synchronised, show how the adjoint method is applied, describe our proposed multi-model setup, and detail the fitting procedure, including the minimisation algorithm. Section 2 shows and discusses the results of our multi-model setups, using a single model setup as a baseline for comparison. The results of introducing a mismodelling term to the adjoint model are also included. Our results are summarised and conclusions discussed in Section 2.3.
1 Methodology
1.1 Lorenz ’63 model
In this study, we use the coupled Lorenz ’63 system as described by Yang et al. (2006). The model is defined by the equations:
$\frac{dx}{dt}$ $=$ $\sigma(y-x),$ (1a)
$\frac{dy}{dt}$ $=$ $\rho x-y-xz,$ (1b)
$\frac{dz}{dt}$ $=$ $xy-\beta z$ (1c)
where $(x,y,z)$ are the state variables at each given time step and $\sigma$, $\rho$, and $\beta$ are the model parameters. Throughout this article, we integrate all our models using the fourth-order Runge-Kutta method with a step size of $\Delta t=0.01$ and total time period of $100$ time units [TUs]. This system of equations will be referred to as the true model for which we will apply the standard values for $\sigma_{t}=10$, $\rho_{t}=28$, and $\beta_{t}=8/3$. This true model is used to generate pseudo-observation which will be used to synchronise our physical models in Sub-section 1.3. Noise is included in these pseudo-observations by adding random values from a Gaussian distribution centred at zero to the true values. The random noise value magnitudes are bounded by a given percentage relative to the systems’ standard distribution. These pseudo-observations will be labelled as $(x_{o},y_{o},z_{o})$.
1.2 Adjoint method
For the Lorenz ’63 system outlined in Eq. (1) the adjoint matrix can be derived by transposing its tangent liner matrix (TLM). The TLM follows from Eq. (1) as:
$M^{T}\equiv\frac{\partial\dot{\vec{x}}_{i}}{\partial\vec{x}_{j}}=\left(\begin{matrix}-\sigma&\sigma&0\\ \rho-z&-1&-x\\ y&x&-\beta\\ \end{matrix}\right).$ (2)
The adjoint $M^{\ast}$ is then given by $M^{\ast}\equiv M^{T}.$ As part of the adjoint model assimilation process, the adjoint equations are integrated backwards in time to calculate the gradient of the cost function with respect to control parameters which are used subsequently in an iterative process to adjust the control parameters such that the system is brought into consistency with observations.
1.3 Synchronisation
A fundamental limitation of the adjoint method arises when integrating over periods that are longer than the predictability time scale of a system, leading to exponentially growing gradients and a cost function with an increasing number of local minima (Köhl and Willebrand, 2002; Lea et al., 2000). The problem can be mitigated by synchronisation which removes the non-linear or chaotic dynamics which effect the cost function (Abarbanel et al., 2010; Sugiura et al., 2014). To incorporate a synchronisation technique, we expand the Lorenz ’63 model by adding so-called nudging terms which then reads:
$\frac{dx_{a}}{dt}$ $=$ $\sigma(y_{a}-x_{a})+\alpha(x_{o}-x_{a}),$ (3a)
$\frac{dy_{a}}{dt}$ $=$ $\rho x_{a}-y_{a}-x_{a}z_{a}+\alpha(y_{o}-y_{a}),$ (3b)
$\frac{dz_{a}}{dt}$ $=$ $x_{a}y_{a}-\beta z_{a}.$ (3c)
Here $\alpha$ is the synchronisation constant, sub-script $a$ denotes the model with an adjoint, $(x_{o},y_{o},z_{o})$ are the pseudo-observations generated from the true model. For this synchronised model we will apply the standard values used in the true system plus a $10\%$-error, which gives $\sigma_{a}=11$, $\rho_{a}=30.8$, and $\beta_{a}=44/15$. These will act as our initial conditions for the parametric fit.
The adjoint matrix for the synchronised model in Eq. (3) is given by:
$M_{a}^{\ast}=\left(\begin{matrix}-(\sigma+\alpha)&\rho-z_{a}&y_{a}\\ \sigma&-(1+\alpha)&x_{a}\\ 0&-x_{a}&-\beta\\ \end{matrix}\right).$ (4)
The synchronisation constant is the only difference from the adjoint of Eq. (2). The significance of this change will be shown in Section 2, as $\alpha$ has a critical role in the precision with which parameters can be estimated, due to its influence on both the cost function and its gradient.
There are seven possible combinations of the three state variables which can be synchronised. The effect of each of the possible choices on the root mean squared error (RMSE) between the true and adjoint systems is shown in Fig. 1 by varying the synchronisation constant $\alpha$ from 0 to 30. The figure demonstrates that synchronising the $z$-component is ineffective at reducing the RMSE (Yang et al., 2006). In contrast, synchronising both $x$ and $y$ prove effective, with $y$ leading the lowest RMSE values of the single variable for all values of $\alpha$. Synchronising $xyz$ and $xy$ achieve the most effective reduction in RMSE for the lowest value of $\alpha$. It can be seen that synchronising $z$ can lead to model instability. Thus, we choose to only synchronise $x$ and $y$ in the following, as seen in Eq. 3, to achieve more stable results with negligible precision loss.
Figure 1: Different coupling schemes are trialled. RMSE is calculated between the truth and model trajectories after synchronisation to pseudo-observations. Noise was added (with zero mean and $\sqrt{2}$ standard deviation) to the truth when constructing the pseudo-observations. The synchronisation constant $\alpha$ is varied from 0 to 30 in steps of 0.5.
The Lorenz ’63 attractors for the trajectories of the true model and that with an adjoint are shown in Fig. 2(a) without synchronisation. A large divergence is visible between the trajectories. However, if synchronisation is introduced, as shown in Fig. 2(b), the trajectories become very similar. There is now significant overlap between their kernel density estimations (KDEs). We choose to use KDEs throughout this paper as they represent a smoothed estimate of the PDF for the trajectory. This allows for convenient visual comparison of trajectories. A more numerically rigorous method to check for effective synchronisation will be discussed in section 2.
((a)) $\alpha=0$.
((b)) $\alpha=10$.
[…]
Figure 2: The bottom left quadrants show the Lorenz ’63 truth and model attractors from the main three variable orientations. The diagonal plots show kernel density estimations (KDEs). Fig. 2(a) shows the trajectories without synchronisation. Fig. 2(b) shows the trajectories with synchronisation.
1.4 Multi-model data assimilation
This technique consecutively synchronises two forward models before running the adjoint of the second model backward in time. For this purpose, Eq. 3 must be modified to incorporate a consecutive synchronisation. A schematic of this setup is provided in Fig.3 and the implications of the two possible ways to calculate the cost function are discussed in the subsequent subsections. The equations of model 1, which is run only in forward mode, are
$\frac{dx_{f}}{dt}$ $=$ $\sigma(y_{f}-x_{f})+\alpha(x_{o}-x_{f}),$ (5a)
$\frac{dy_{f}}{dt}$ $=$ $\rho x_{f}-y_{f}-x_{f}z_{f}+\alpha(y_{o}-y_{f}),$ (5b)
$\frac{dz_{f}}{dt}$ $=$ $x_{f}y_{f}-\beta z_{f}$ (5c)
where the sub-script $f$ denotes the forward run of model 1 and sub-script $o$ denotes observations generated from truth model. The system of equations for the model 2 which has an adjoint will now be modified to synchronise with the forward-only model and not the observations:
$\frac{dx_{a}}{dt}$ $=$ $\sigma(y_{a}-x_{a})+\alpha(x_{f}-x_{a}),$ (6a)
$\frac{dy_{a}}{dt}$ $=$ $\rho x_{a}-y_{a}-x_{a}z_{a}+\alpha(y_{f}-y_{a}),$ (6b)
$\frac{dz_{a}}{dt}$ $=$ $x_{a}y_{a}-\beta z_{a}$ (6c)
where the sub-script $a$ denotes model 2 which has an adjoint. This model synchronises with the model 1 but never directly with the observations.
Figure 3: Illustration of the multi-model setup where each pseudo-observation generated from the truth includes random additive Gaussian noise. The cost function can measure the difference between the observations and either model 1 or 2 depending on the assumptions made. Both options are discussed in the text.
To assimilate the data, we fit one of the synchronised models to the observations by optimising the model parameters. A cost function is constructed to calculate the misfit between observations and the model of interest. The gradient of the cost function, with respect to the model parameters, is always calculated using the adjoint method associated with model 2. However, the form of the adjoint will vary between the two methods we subsequently present. The adjoint model is numerically evaluated by automatic differentiation (AD) of the forward model 2 with respect to either model 1 or 2 depending on our chosen setup. This is done in the python package JAX which numerically evaluates the vector Jacobian product of the model with respect to its state variable vector (Bradbury et al., 2018). This is then integrated using an inverse Runge-Kutta scheme. Our code stores the state variables and adjoint vectors at each time step. It is also possible to carry out the entire integration using JAX. The process of synchronising all models, calculating the cost function and its gradient, and then adjusting model parameters is carried out iteratively by our chosen minimisation algorithm. Throughout all steps the parameter value of forward-only and adjoint models are identical and optimised simultaneously.
[…]
1.4.1 Setup 1 - state filtered data assimilation (SFDA)
Assuming that both Lorenz models can be thought of as representing two identical climate models, the cost function can be placed on the model 2. This allows model 1 to filter out some of the background noise on the observations before they are given to the cost function attached to model 2. This favourably increases the signal to background ratio synchronised into model 2. Such a filtering setup would theoretically reduce parametric uncertainty below that of traditional single model data assimilation because model 1 should act to reduce the amount of noise synchronised into model 2. We will subsequently refer to this setup as state filtered data assimilation (SFDA.)
In SFDA the cost function acts to constrain model 2. The cost function is
$J_{\text{SFDA}}=\frac{1}{2N}\sum_{t=0}^{N}\left(\vec{x}_{o}(t)-\vec{x}_{a}(t)\right)^{T}\frac{1}{\sigma^{2}_{\vec{x}_{o}}}\left(\vec{x}_{o}(t)-\vec{x}_{a}(t)\right)$ (7)
$N$ is the total number of time steps of the assimilation window, and $\sigma_{\vec{x}_{o}}$ is the uncertainty associated with the observation noise. The adjoint matrix includes additional terms, compared with Eq. 4, arising from the second synchronisation step for model 2. This is given by:
$M_{\text{SFDA}}^{\ast}=\left(\begin{matrix}-(\sigma+\alpha)&\rho-z_{a}&y_{a}&0&0&0\\ \sigma&-(1+\alpha)&x_{a}&0&0&0\\ 0&-x_{a}&-\beta&0&0&0\\ \alpha&0&0&-(\sigma+\alpha)&\rho-z_{f}&y_{f}\\ 0&\alpha&0&\sigma&-(1+\alpha)&x_{f}\\ 0&0&0&0&-x_{f}&-\beta\\ \end{matrix}\right)$ (8)
which in practice is numerically evaluated using AD.
1.4.2 Setup 2 - hybrid data assimilation (HDA)
Here we want to explore if using an existing adjoint from one model could be utilised to optimise a different target model without adjoint. This will be referred to as hybrid data assimilation (HDA.) In HDA we assume that both models may differ in resolution or numerical formulation but are governed by the same equations. Instead of interpolating or transforming the original model variables onto the adjoint model grid, formulation of the adjoint model through synchronisation would provide a simpler means to do this, as only essential parameters need to be interpolated. Auxiliary variables and parameters, including those that may not exist in the target model, will be generated by the synchronised model.
The cost function of HDA is
$J_{\text{HDA}}=\frac{1}{2N}\sum_{t=0}^{N}\left(\vec{x}_{o}(t)-\vec{x}_{f}(t)\right)^{T}\frac{1}{\sigma^{2}_{\vec{x}_{o}}}\left(\vec{x}_{o}(t)-\vec{x}_{f}(t)\right).$ (9)
$N$ is the total number of time steps of the assimilation window and $\sigma_{\vec{x}_{o}}$ is the uncertainty associated with the observation noise. This measures the quadratic misfit between the forward-only model 1 and the observations. Model 1 $\vec{x}_{f}$ will be constrained by this cost function and its gradient will be calculated using the adjoint of model 2 $\vec{x}_{a}$. The adjoint matrix is
$M_{\text{HDA}}^{\ast}=\left(\begin{matrix}-(\sigma+\alpha)&\rho-z_{a}&y_{a}\\ \sigma&-(1+\alpha)&x_{a}\\ 0&-x_{a}&-\beta\\ \end{matrix}\right).$ (10)
which is numerically evaluated using AD. The synchronisation with model 1 ensures that the trajectory of model 2 $\vec{x}_{a}$ closely follows that of model 1.
1.4.3 Cost function gradient
The gradient of the cost function, with respect to the state variables at $t=0$, is given by
$\vec{\nabla}_{\vec{x}_{a}}J=\frac{1}{N}\vec{\lambda}(0),$ (11)
where $\vec{\lambda}$ is the adjoint vector. The adjoint vector is calculated by integration of the differential adjoint equation in the reverse time direction. This will differ between our two setups.
SFDA (setup 1)
This adjoint equation for SFDA is given by:
$\dot{\vec{\lambda}}_{\text{SFDA}}(t)$ $=$ $\frac{1}{\sigma_{\vec{x}_{o}}^{2}}\left(\vec{x}_{o}(t)-\vec{x}_{a}(t)\right)$
$-M_{\text{SFDA}}^{\ast}(t)\vec{\lambda}_{\text{SFDA}}(t)\leavevmode\nobreak\ \text{for}\leavevmode\nobreak\ t=N,...,0$
$\text{with}\leavevmode\nobreak\ \leavevmode\nobreak\ \vec{\lambda}_{\text{SFDA}}(N)$ $=$ $\vec{0}.$ (12b)
These equations were derived using the method detailed in Talagrand (2010). The gradient can then be calculated with respect to the parameters $(\sigma,\rho,\beta)$ notated by the subscript $\vec{\theta}$. This yields
$\vec{\nabla}_{\vec{\theta}}J_{\text{SFDA}}=\frac{1}{N}\sum_{t=N}^{0}\vec{\lambda}_{\text{SFDA}}(t)\left(\begin{matrix}y_{a}(t)-x_{a}(t)\\ x_{a}(t)\\ -z_{a}(t)\\ y_{f}(t)-x_{f}(t)\\ x_{f}(t)\\ -z_{f}(t)\end{matrix}\right)$ (13)
which is a component-wise multiplication at each time step.
HDA (setup 2)
This adjoint equation for HDA is given by:
$\dot{\vec{\lambda}}_{\text{HDA}}(t)$ $=$ $\frac{1}{\sigma_{\vec{x}_{o}}^{2}}\left(\vec{x}_{o}(t)-\vec{x}_{f}(t)\right)$
$-M_{\text{HDA}}^{\ast}(t)\vec{\lambda}_{\text{HDA}}(t)\leavevmode\nobreak\ \text{for}\leavevmode\nobreak\ t=N,...,0$
$\text{with}\leavevmode\nobreak\ \leavevmode\nobreak\ \vec{\lambda}_{\text{HDA}}(N)$ $=$ $\vec{0}.$ (14b)
The gradient with respect to the parameters $\vec{\theta}=(\sigma,\rho,\beta)$ is calculated to be:
$\vec{\nabla}_{\vec{\theta}}J_{\text{HDA}}=\frac{1}{N}\sum_{t=N}^{0}\vec{\lambda}_{\text{HDA}}(t)\left(\begin{matrix}y_{a}(t)-x_{a}(t)\\ x_{a}(t)\\ -z_{a}(t)\end{matrix}\right)$ (15)
which is again a component-wise multiplication at each time step. For HDA and SFDA, the trajectories of both models and adjoint vectors are stored for evaluation of the gradient.
2 Results
Throughout the following section we will use the single model described in Lyu et al. (2018) as our benchmark to compare the new setups against. To understand the behaviour of the setups at different operating extremes, assimilations are carried out for variations of observational noise and $\alpha$. This will help establish the optimal synchronisation strength dependent on the noise amplitude. We will also be able to compare the errors and uncertainties of the single model with our multi-model setups.
[…]
2.1 SFDA (setup 1)
The results from a scan of $\alpha$ are shown in Fig. 4. The single model scan has two main regions. The first, $\alpha\leq 7.5$, is where the system is poorly synchronised giving an inaccurate fit of the parameters. The second, above $\alpha>7.5$, is where the system is fully synchronised and recovers the true model parameters very effectively. SFDA has a higher onset of effective synchronisation than the single model setup, beginning at $\alpha=11$. Above $\alpha=12.5$, SFDA has consistently more accurate parameter recovery than the single model setup.
Figure 4: The percentage error between the true values of $(\sigma,\rho,\beta)$ and the fitted value from SFDA. A single model assimilation is included for comparison. An ensemble of 100 assimilations is carried out over 100 different data sets. The median (lines) and $68\%$ percentile intervals (shaded areas) are plotted. The noise level is $25\%$.
Fig. 5 shows the results of two fits carried out with noise of $25\%$. The mean percentage uncertainty over the three parameters is plotted for both setups. Noticing in particular the spread of the percentile intervals, the single model setup is found to be synchronised and have a high precision from $\alpha=7.5$ and SFDA from $\alpha=11.5$. Once the SFDA setup is synchronised, it is found to have a reduced uncertainty compared with the single model. SFDA is found to be approximately one third more precise better than the single model setup for all value of $\alpha$ investigated.
Fig. 6 shows the results of varying the noise levels on the fitted parameter values. For all noise levels studied the quality of the fit can be considered good as the median of the mean percentage uncertainty on the parameters remains below 0.5% even with noise levels of up to 50%. SFDA is found to have similar mean error performance to the single model system across the range of noise levels tested. The spread of the error is slightly improved in the double model setup at low noise likely due to the forward-only model ‘smoothing’ outlying observation better than a single model setup. The parametric uncertainty is found to be consistently reduced in the double model system for all noise levels. This demonstrates the precision improvement achieved by running the forward model twice to ‘smooth’ the observations before carrying out data assimilation. The consequences of this are that for smaller or localised climate models, where computational resources are available and improved precision is desirable, SFDA could reduce error and particularly uncertainty.
Figure 5: The average percentage uncertainty on the three parameters $(\sigma,\rho,\beta)$ after SFDA from the minimisation algorithm for different values of $\alpha$. A single model assimilation is included for comparison. An ensemble of 100 assimilations is carried out over 100 different data sets. The median (line) and $68\%$ percentile intervals (shaded areas) are plotted.
((a)) Percentage error.
((b)) Percentage uncertainty.
Figure 6: The percentage error between the true values of $(\sigma,\rho,\beta)$ and those from SFDA, as well as average percentage uncertainty on the SFDA parameters. A single model assimilation is included for comparison. An ensemble of 100 assimilations is carried out over 100 different data sets. The median (line) and $68\%$ percentile intervals (shaded areas) are plotted. The noise level varies between 5% and 50% in steps of 5%.
2.2 HDA (setup 2)
The results from a scan of $\alpha$ are shown in Fig. 7. The single model scan has two main regions. The first, $\alpha\leq 7.5$, is where the system is poorly synchronised giving an inaccurate fit of the parameters. The second, $\alpha\geq 7.5$, is where the system is synchronising very effectively and recovers the true model parameters very accurately. The HDA scan follows the behaviour of the primary model very closely.
Figure 7: The percentage error between the true values of $(\sigma,\rho,\beta)$ and those from HDA. A single model assimilation is included for comparison. An ensemble of 100 assimilations is carried out over 100 different pseudo-data sets. The median (lines) and $68\%$ percentile intervals (shaded areas) are plotted. The noise level is $25\%$.
Fig. 8 shows the results of two fits carried out with noise of $25\%$. The mean percentage uncertainty over the three parameters is plotted for both setups. HDA is found to have almost identical uncertainty to the single model. The plot consists of two regions. The first, for $\alpha<7.5$ is the region where the model is not yet consistently synchronised producing high variability depending on the specific noise. The second, $\alpha\geq 7.5$ is where the system is consistently synchronised. The minimum median of the mean parametric uncertainty, after consistent synchronisation begins, is $\approx 0.35\%$ and achieved at $\alpha=7.5$. The subsequent increase in uncertainty is due to the increased $\alpha$ both reducing and flattening the cost function which reduces the parametric sensitivity of the fit.
Figure 8: The percentage uncertainty, from the minimisation algorithm, averaged over all three parameters $(\sigma,\rho,\beta)$ after HDA. A single model assimilation is included for comparison. An ensemble of 100 assimilations is carried out over 100 different data sets. The median (lines) and $68\%$ percentile intervals (shaded areas) are plotted.
[…]
((a)) Percentage error.
((b)) Percentage uncertainty.
Figure 9: The percentage error between the true values of $(\sigma,\rho,\beta)$ and those from HDA. A single model assimilation is included for comparison. An ensemble of 100 assimilations is carried out over 100 different data sets. The median (lines) and $68\%$ percentile intervals (shaded areas) are plotted. The noise level varies between 5% and 50% in steps of 5%.
2.3 Mismodelling in HDA
HDA (setup 2) uses different forward and adjoint models with some shared physics. Therefore, it is important to examine the impact of a difference in the physics of these two models. We construct a test case where the equations of model 2, which has an adjoint, Eq. (6) are modified to give an oscillatory difference to both the truth and model 1. This can be done in a number of ways. We chose to introduce a multiplicative sine function to the equations in such a way it is also included in the adjoint matrix and thus modifies the gradient values returned to the fitting algorithm. Model 2 with an adjoint is now
$\frac{dx_{a}}{dt}$ $=$ $\sigma(y_{a}-x_{a})+\alpha(x_{f}-x_{a}),$ (16a)
$\frac{dy_{a}}{dt}$ $=$ $\rho x_{a}-y_{a}-x_{a}z_{a}+\alpha(y_{f}-y_{a}),$ (16b)
$\frac{dz_{a}}{dt}$ $=$ $x_{a}y_{a}-\beta z_{a}\cdot\left(1-\epsilon\sin\left(2\pi t\right)\right)$ (16c)
where $\epsilon$ is term which determines the strength of the oscillation term and $t$ is the simulation time. The effect of this term on the attractor is shown in Fig. 10, without synchronisation and assimilation. When compared to Fig. 2(a), it can be seen that this term is very effective at distorting the shape and probability density of the attractor.
[…]
Figure 10: The Lorenz truth and model 2 attractors in the case of $\epsilon=1.0$. The bottom left and top right quadrants shows the attractors from all possible co-ordinate orientations. The diagonal plots show kernel density estimations (KDEs). No noise is added.
The consequence of varying this term on the accuracy of parameter optimisation after assimilation are shown in Fig. 11. With increasing epsilon the percentage error and uncertainty between fitted and truth systems remains stable. In spite of the large impact this term has on the attractor shape, this demonstrates a resilience of HDA to modelling differences between the forward-only model 1 and model 2 with an adjoint.
Figure 11: The percentage error (left) and uncertainty (right) between the true values of $(\sigma,\rho,\beta)$ and those from HDA. An ensemble of 100 assimilations is carried out over 100 different data sets. The median (line) and $68\%$ percentile intervals (shaded areas) are plotted. The noise level $25\%$ and, $\alpha=7.5$.
[…]
In this paper we have demonstrated the ability to constrain a Lorenz ’63 model using a second model with an adjoint by 4D-Var data assimilation. Such an approach removes the need to generate an adjoint for a forward model, if such an adjoint already exists for a separate, yet similar system. An important application of this technique in Earth system modelling would be a situation where a low-resolution ESM with an adjoint shares a parametrisation with a high-resolution ESM without an adjoint. The low-resolution ESM could, through synchronisation, provide all necessary variables to run its tangent linear adjoint model. This can then be utilised to estimate parameters in the high-resolution ESM without an adjoint. Moreover, using a simpler lower resolution version could make data assimilation quicker and would use fewer computational resources than an adjoint of the high-resolution ESM. We have also shown that running a forward model twice before beginning data assimilation can act to smooth the data and reduce the parametric uncertainty by roughly one third compared to a single forward run. Our overarching attention in this approach is optimising the parameters of a full ESM which will be tested as a next step. Nevertheless, it would also be possible to optimise the state variables which are more applicable to weather forecasting techniques. Future work will examine the resilience of such setups to spacially and temporally sparse data.
[…]
The pseudo-data samples used in this study are available on request from the authors.
[…]
PDK carried out the research, wrote the initial manuscript draft, and edited it. AB also assisted with the research and edited the manuscript. AK and DS designed the research concept and edited the manuscript.
[…]
The authors declare no potential competing interest.
Acknowledgements.
The figures in this research were plotted by the seaborn plotting package (Waskom, 2021) utilising our output which was managed using pandas (pandas development team, 2020). Automatic differentiation of our model was carried out using JAX (Bradbury et al., 2018). The minimisation of our cost function and uncertainty evaluation was done by iminuit (Dembinski and et al., 2020). This research was supported, in part, through the Koselleck grant $\textit{Earth}^{\textit{RA}}$ from the Deutsche Forschungsgemeinschaft (grant number: 66492934.) Contribution to the Centrum für Erdsystemforschung und Nachhaltigkeit (CEN) of Universität Hamburg. This work used resources of the Deutsches Klimarechenzentrum (DKRZ) granted by its Scientific Steering Committee (WLA).